OpenStudy (anonymous):

There are 17 broken light bulbs in a box of 100 light bulbs. A random sample of 3 light bulbs is chosen without replacement. How many light bulbs contain exactly exactly 1 broken light bulb?

OpenStudy (tkhunny):

Please expand: \((p+q)^{3}\) Trust me on this.

OpenStudy (anonymous):

p^3+3qp^2+3pq^2+q^3

OpenStudy (tkhunny):

That's pretty good. p = Probability that a light is working q = Probability that a light is broken = 1-p Can you now solve the problem?

OpenStudy (anonymous):

No >.< . Besides, I was asked to use combinations/Permutations .

OpenStudy (tkhunny):

Why didn't you say so? You need to: Select 2 of 83 good. Select 1 of 17 bad. from a selection of 3 from 100.

OpenStudy (anonymous):

Sorry :P .

OpenStudy (anonymous):

Yeah that's what's confusing me. Why is it 2 from 83? 83 are the bulbs that are NOT broken.

OpenStudy (tkhunny):

1 bad means 2 good to go with the bad one.

OpenStudy (anonymous):

Ahh makes sense!

OpenStudy (anonymous):

So really it would be (83 C 2)*(17 C 1)

OpenStudy (anonymous):

If i'm correct.

OpenStudy (anonymous):

Yep. That is indeed correct. Thanks a lot!

OpenStudy (anonymous):

That's actually incorrect :P . . That would be how many samples contain no broken light bulbs.

OpenStudy (anonymous):

83*82*81/100*99*98=30627/53900

OpenStudy (anonymous):

one is defective so (1-30627/53900)

OpenStudy (anonymous):

23273/53900

OpenStudy (tkhunny):

The binomial approximation that I tried to get you to do up front suggests the correct probability is in the neighborhood of \(3\cdot0.83^{2}\cdot 0.17 = 0.351\) You should try: \(\dfrac{ _{83} C_{2}*_{17} C_{1} }{_{100} C_{3}}\)

OpenStudy (anonymous):

0.4317 is it correct

OpenStudy (tkhunny):

No. Please be more systematic and make an attempt to explain what it is you are doing.

OpenStudy (anonymous):

p(none is defective)=83c3/100c3=30627/53900 p(one is defective)=1-30627/53900

OpenStudy (tkhunny):

How about p(2 are defective) and p(3 are defective)?

OpenStudy (anonymous):

exactly exactly 1 broken light bulb?

OpenStudy (tkhunny):

No, p(none broken) + p(1 broken) is NOT unity (1). No, p(none broken) + p(ANY broken) is unity (1). This is what you calculated, ANY, not ONE.

OpenStudy (tkhunny):

p(0 broken) + p(1 broken) + p(2 broken) + p(3 broken) = p(0 broken) + p(any broken)

OpenStudy (tkhunny):

Please be more systematic and make an attempt to explain what it is you are doing.

OpenStudy (tkhunny):

No. That is now farther off than it was before. Look a few posts up. The answer is given correctly as (83C2)(17C1)/(100C3)

OpenStudy (anonymous):

p one is defective=17/100*83/99*82/98

OpenStudy (tkhunny):

Okay, now you're just writing things down. Please stop doing it. Go look at the correct answer and learn from it. Stop trying to guess that it might be something else. In this case, you have entirely forgotten that the three can be drawn in any order.

OpenStudy (tkhunny):

Really? Then why does it give an incorrect answer? Take a very hard look at the binomial approximation. p^3+3qp^2+3pq^2+q^3 This would be exact for WITH replacement. Since there are so few drawn, from such a large population, this is very close and very clear. In particular, why are those middle coefficients "3" instead of "1"?

OpenStudy (anonymous):

but u should add the answer i think p(none is defective)=83c3/100c3=30627/53900 +P(one defective)=17/100*83/99*82/98+83/100*82/99*17/98+83/100*17/99*82/98

OpenStudy (anonymous):

i.e 0.568+0.378=0.9462

OpenStudy (tkhunny):

And 1-0.9462 = p(2) + p(3)